# RCDN: Towards Robust Camera-Insensitivity Collaborative Perception via Dynamic Feature-based 3D Neural Modeling

### v1

@@ Abstract
Collaborative perception is dedicated to tackling the constraints of single-agent perception, such as occlusions, based on the multiple agents’ multi-view sensor inputs. However, most existing works assume an ideal condition that all agents’ multi-view cameras are continuously available. In reality, cameras may be highly noisy, obscured or even failed during the collaboration. In this work, we introduce a new robust camera-insensitivity problem: how to overcome the issues caused by the failed camera perspectives, while stabilizing high collaborative performance with low calibration cost? To address above problems, we propose RCDN, a Robust Camera-insensitivity collaborative perception with a novel Dynamic feature-based 3D Neural modeling mechanism. The key intuition of RCDN is to construct collaborative neural rendering field representations to recover failed perceptual messages sent by multiple agents. To better model collaborative neural rendering field, RCDN first establishes a geometry BEV feature based time-invariant static field with other agents via fast hash grid modeling. Based on the static background field, the proposed time-varying dynamic field can model corresponding motion vectors for foregrounds with appropriate positions. To validate RCDN, we create OPV2V-N, a new large-scale dataset with manual labelling under different camera failed scenarios. Extensive experiments conducted on OPV2V-N show that RCDN can be ported to other baselines and improve their robustness in extreme camera-insensitivity settings. Our code and datasets will be available soon.
@@ 1 Introduction
To address this robust camera insensitivity collaborative perception problem, we propose RCDN, a Robust Camera-insensitivity collaborative perception with a Dynamic feature-based 3D Neural modeling mechanism. The core idea is to recover noisy camera perceptual information from other agents’ views by modeling the collaborative neural rendering field representations. Specifically, RCDN has two collaborative field phases: a time-invariant static background field and time-varying dynamic foreground field. In the static phases, RCDN sets other baselines’ backbone as the collaboration base and undertakes end-to-end training to create a robust unified geometry BEV feature space for all agents. Then, the geometry BEV feature combines the hash grid modeling, an explicit and multi-resolution network, to generate static background views through $\alpha$-composed accumulation of RGB values along a ray at a fast speed. In the dynamic phase, RCDN utilizes 4D spatiotemporal position features to model the dynamic motion of 3D points, which learns an accurate motion field under optical priors and spatiotemporal regularization. The proposed RCDN has two major advantages: i) RCDN can handle camera insensitivity collaboration under unknown noisy timestamps and numbers; ii) RCDN does not put any extra communication burden into inference stage and costs little computation burden.
@@ 4.4 Dynamic Collaborative Neural Field
where $G_{\theta}^{d}$ shares the same hash grid representations, but for the dynamic collaborative neural field optimization; $\Delta(\cdot,\cdot)$ is the temporal interpolation functions, which makes the $\mathrm{MLP}$ can efficiently learn the features between keyframes in a scalable manner. Meanwhile, to improve the temporal consistency of the proposed field, we compute the collaborative scene flow neighbors $\mathbf{r}(u_{k})+\mathbf{s}_{fw}$ and $\mathbf{r}(u_{k})-\mathbf{s}_{bw}$ with the predicted collaborative scene flow $\mathbf{s}_{fw},\mathbf{s}_{bw}$ to warp the collaborative neural field from the neighboring time instance to the current time. Hence, we can obtain the corresponding density and color of adjacent time by querying the same $\mathrm{MLPs}$ model at $\mathbf{r}(u_{k})+\mathbf{s}$:
@@ 4.5 Training Details and Optimization
To train the overall system, we supervise two tasks: static and dynamic collaborative neural fields, respectively. Meanwhile, during the training process, the static collaborative field and dynamic collaborative field are trained separately. The total reconstruction loss is a weighted combination as the following:
@@ 5.4 Ablation Study
We conduct ablation studies on OPV2V-N with the CoBEVT baseline method. Table 4 assesses the effectiveness of the proposed two field phases. We see that i) only one neural field can recover most static part performance from the noisy camera data; ii) the proposed time model in collaborative dynamic fields can handle the motion blurry caused by the vehicles, shown in Figure 4. Meanwhile, we compare the training efficiency of the proposed RCDN with existing dynamic fields modeling methods[68], as illustrated in Figure 6. Our approach, which leverages explicit grid and geometry feature-based representations, accelerates the training process by approximately 24×24$\times$24 × compared to the existing implicit MLP-based modeling, while also achieving superior PSNR quality. See more discussions in the Appendix B.2.
@@ Appendix B Detailed Information about Experiments
B.3 Benchmarks
@@ B.4 PSNR Results
B.5 Geometry BEV Volume Feature
@@ Appendix C Visualization
We visualize some segmentation results of different baseline methods w.o/w RCND under the different scenes, shown in Figure 14-18.

### v2

@@ Abstract
Collaborative perception is dedicated to tackling the constraints of single-agent perception, such as occlusions, based on the multiple agents’ multi-view sensor inputs. However, most existing works assume an ideal condition that all agents’ multi-view cameras are continuously available. In reality, cameras may be highly noisy, obscured or even failed during the collaboration. In this work, we introduce a new robust camera-insensitivity problem: how to overcome the issues caused by the failed camera perspectives, while stabilizing high collaborative performance with low calibration cost? To address above problems, we propose RCDN, a Robust Camera-insensitivity collaborative perception with a novel Dynamic feature-based 3D Neural modeling mechanism. The key intuition of RCDN is to construct collaborative neural rendering field representations to recover failed perceptual messages sent by multiple agents. To better model collaborative neural rendering field, RCDN first establishes a geometry BEV feature based time-invariant static field with other agents via fast hash grid modeling. Based on the static background field, the proposed time-varying dynamic field can model corresponding motion vectors for foregrounds with appropriate positions. To validate RCDN, we create OPV2V-N, a new large-scale dataset with manual labelling under different camera failed scenarios. Extensive experiments conducted on OPV2V-N show that RCDN can be ported to other baselines and improve their robustness in extreme camera-insensitivity settings.
@@ 1 Introduction
To address this robust camera insensitivity collaborative perception problem, we propose RCDN, a Robust Camera-insensitivity collaborative perception with a Dynamic feature-based 3D Neural modeling mechanism. The core idea is to recover noisy camera perceptual information from other agents’ views by modeling the collaborative neural rendering field representations. Specifically, RCDN has two collaborative field phases: a time-invariant static background field and time-varying dynamic foreground field. In the static phases, RCDN sets other baselines’ backbone as the collaboration base and undertakes end-to-end training to create a robust unified geometry Bird-eye view (BEV[28, 29]) feature space for all agents. Then, the geometry BEV feature combines the hash grid modeling, an explicit and multi-resolution network, to generate static background views through $\alpha$-composed accumulation of RGB values along a ray at a fast speed. In the dynamic phase, RCDN utilizes 4D spatiotemporal position features to model the dynamic motion of 3D points, which learns an accurate motion field under optical priors and spatiotemporal regularization. The proposed RCDN has two major advantages: i) RCDN can handle camera insensitivity collaboration under unknown noisy timestamps and numbers; ii) RCDN does not put any extra communication burden into inference stage and costs little computation burden.
@@ 4.4 Dynamic Collaborative Neural Field
where $G_{\theta}^{d}$ shares the same hash grid representations, but for the dynamic collaborative neural field optimization; $\Delta(\cdot,\cdot)$ is the temporal interpolation functions, which makes the $\mathrm{MLP}$ can efficiently learn the features between keyframes in a scalable manner. Meanwhile, to improve the temporal consistency of the proposed field, we compute the collaborative scene flow neighbors $\mathbf{r}(u_{k})+\mathbf{s}_{fw}$ and $\mathbf{r}(u_{k})-\mathbf{s}_{bw}$ with the predicted collaborative scene flow $\mathbf{s}_{fw},\mathbf{s}_{bw}$ to warp the collaborative neural field from the neighboring time instance to the current time. Note that the term $\mathbf{s}_{fw}$ stands for forward scene flow, while $\mathbf{s}_{bw}$ refers to backward scene flow. Specifically, the forward scene flow ($\mathbf{s}_{fw}$) estimates the flow from time t to t+1, whereas the backward scene flow ($\mathbf{s}_{bw}$) estimates the flow from time t to t-1. Hence, we can obtain the corresponding density and color of adjacent time by querying the same $\mathrm{MLPs}$ model at $\mathbf{r}(u_{k})+\mathbf{s}$:
@@ 4.5 Training Details and Optimization
To train the overall system, we supervise two tasks: static and dynamic collaborative neural fields, respectively. Meanwhile, during the training process, the static collaborative field and dynamic collaborative field are trained separately. The initial learning rate is 5e-4 with the exponential learning rate decay strategy. The weight values are set to 1.0, 1.0, 0.1, and 1.0, respectively:
@@ 5.4 Ablation Study
Components analysis We conduct ablation studies on OPV2V-N with the CoBEVT baseline method. Table 4 assesses the effectiveness of the proposed two field phases. We see that i) only one neural field can recover most static part performance from the noisy camera data; ii) the proposed time model in collaborative dynamic fields can handle the motion blurry caused by the vehicles, shown in Figure 4. Meanwhile, we compare the training efficiency of the proposed RCDN with existing dynamic fields modeling methods[70], as illustrated in Figure 6. Our approach, which leverages explicit grid and geometry feature-based representations, accelerates the training process by approximately 24×24$\times$24 × compared to the existing implicit MLP-based modeling, while also achieving superior PSNR quality. See more discussions in the Appendix B.2.
Performance bottlenecks Regarding the increasing number of agents and cameras, we validated the impact of adding more cameras using the OPV2V-N dataset (corresponding scenario types are T section and midblock respectively) with the CoBEVT baseline. From Table 3, we observe the following: i) With a single overlapping camera view, the proposed method significantly improves baseline performance, and ii) While theoretically, more cameras can provide a larger overlap range, the addition of multiple cameras (depending on their positions) may introduce redundant viewing angles, resulting in less significant performance improvements.
Table 3: Map-view segmentation performance validation about the increasing number of cameras under OPV2V-N datasets with CoBEVT baseline. Note that the failure setting is under one random noisy camera failure in the testing phases. We report IoU for all classes.
Methods Metrics Scene Failure Overlap Cameras
+1 +2 +3
CoBEVT[6] Dr. Area T Section 23.23 26.97 26.91 27.23
Midblock 23.43 38.87 38.94 39.51
Dyn. Vehicles T Section 18.83 40.72 41.38 42.29
Midblock 16.57 45.60 48.31 49.88
Figure 7: Visualization of proposed RCDN for detection downstream task performance. Note that red and green boxes denote detection results and ground-truth respectively.
Table 4: Detection performance of CoBEVT and V2VNet baseline methods w.o/w. the proposed RCDN on OPV2V-N dataset with one random noisy camera failure in the testing phase. We report Average Precision (AP) at Intersection-over-Union (IoU) thresholds of 0.50 and 0.70.
Methods / Metrics Normal
AP@0.50($\uparrow$) AP@0.70($\uparrow$)
CoBEVT[6] 55.56 43.33
V2VNet[37] 58.77 42.42
Methods / Metrics Failures
w.o/w. RCDN
AP@0.50($\uparrow$) AP@0.70($\uparrow$)
CoBEVT[6] 46.67/55.56 34.57/43.21
V2VNet[37] 45.45/53.85 36.37/38.18
Different downstream tasks Our proposed RCDN is general to different downstream tasks and is not limited to just BEV segmentation. We focus on BEV segmentation due to its crucial role in autonomous driving, with direct applications to other tasks such as layout mapping, action prediction, route planning, and collision avoidance. Additionally, we have validated RCDN for detection tasks, shown in Figure 7. We replaced the original segmentation header with a detection header in our experiments. Table 7 shows that for CoBEVT, using RCDN improves the metrics of AP@0.50 and AP@0.70 by 19.05% and 24.99%, respectively.
@@ 6 Conclusion and Limitation
7 Acknowledgments
This work was supported by the National Key Research and Development Program of China (No. 2021YFB2501104), in part by the National Natural Science Foundation of China (No. 62372329), in part by Shanghai Scientific Innovation Foundation (No. 23DZ1203400), in part by Tongji-Qomolo Autonomous Driving Commercial Vehicle Joint Lab Project, and in part by Xiaomi Young Talents Program.
@@ Appendix B Detailed Information about Experiments
B.3 Multi-agents Collaborative Perception
The MCP module stands for the Multi-agent Collaborative Perception process. Existing state-of-the-art (SoTA) MCP modules share a common pipeline: an encoder-fusion-decoder architecture. To ensure fairness in collaborative perception experiments, different MCP modules use the same encoder-decoder architecture but differ in the fusion process. The fusion process is responsible for the bird-eye view (BEV) feature aggregation. Therefore, the MCP module can be replaced by simply switching between different BEV feature aggregation processes.
B.4 Benchmarks
@@ B.5 PSNR Results
B.6 Geometry BEV Volume Feature
@@ Appendix C Visualization
We visualize some segmentation results of different baseline methods w.o/w RCND under the different scenes, shown in Figure 15-19.